OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

lol i knew that haha jk thanks so much! :D

OpenStudy (ahsome):

Let me just put the steps for others to see :) Equation: \[3(x + 1) = -2(x - 1) - 4.\] Expand brackets\[3x+3=-2x+2-4\] Simplify RHS \[3x+3=-2x-2\] Move the \(3\) to the RHS, move the \(-2x\) to the LHS \[3x+2x=-2-3\]\[5x=-5\] Divide both sides by \(5\) \[x=\frac{-5}{5}\]\[x=1\]

OpenStudy (ahsome):

Just so that others can see @horselover33